# SuperARC: A Test for General and Super Intelligence Based on First Principles of Recursion Theory and Algorithmic Probability

### v1

@@ Abstract
We introduce an open-ended test grounded in algorithmic probability that can avoid benchmark contamination in the quantitative evaluation of frontier models in the context of their Artificial General Intelligence (AGI) and Superintelligence (ASI) claims. Unlike other tests, this test does not rely on statistical compression methods (such as GZIP or LZW), which are more closely related to Shannon entropy than to Kolmogorov complexity. The test challenges aspects related to features of intelligence of fundamental nature such as synthesis and model creation in the context of inverse problems (generating new knowledge from observation). We argue that metrics based on model abstraction and optimal Bayesian inference for planning can provide a robust framework for testing intelligence, including natural intelligence (human and animal), narrow AI, AGI, and ASI. Our results show no clear evidence of LLM convergence towards a defined level of intelligence, particularly AGI or ASI. We found that LLM model versions tend to be fragile and incremental, as new versions may perform worse than older ones, with progress largely driven by the size of training data. The results were compared with a hybrid neurosymbolic approach that theoretically guarantees model convergence from optimal inference based on the principles of algorithmic probability and Kolmogorov complexity. The method outperforms LLMs in a proof-of-concept on short binary sequences. Our findings confirm suspicions regarding the fundamental limitations of LLMs, exposing them as systems optimised for the perception of mastery over human language. Progress among different LLM versions from the same developers was found to be inconsistent and limited, particularly in the absence of a solid symbolic counterpart.
Keywords: Abstraction and Reasoning Corpus (ARC), Artificial General Intelligence, prediction, compression, program synthesis, inverse problems, symbolic regression, comprehension, Superintelligence, Generative AI, symbolic computation, hybrid computation, Neurosymbolic computation.
@@ 1 Introduction
We are heavily biased to believe that the way we think and act represents the acme of intelligence, even in instances where we may be limited, or flawed or irrational, or engaged in narrowly specific human (and often mundane) activities like chatting or washing dishes.
@@
One of the greatest lessons and realisations from the impressive apparent performance of Large Language Models (LLMs) is that language and potentially other areas of human intellect are overrated and are more reliant than we thought on memorisation and statistical pattern-matching therefore indicating that these features are not a good benchmark for machine intelligence in the context of features we use to attribute greater value, such as model abstraction and predictive planning.
One of the first metrics for intelligence was introduced by Charles Spearman in 1904 [1]. He proposed specific tests called ‘s’ that would each contribute to a general intelligence test under the name ‘g’, representing the common cognitive ability underlying performance in various mental tasks. Specific intelligences contributing to the estimation of the g factor are verbal comprehension, perceptual reasoning, working memory, processing speed, quantitative reasoning, abstract reasoning, spatial ability, memory retrieval, auditory processing, and fluid reasoning. Some LLM benchmarks test for different factors, with several benchmarks based on correct answers versus hallucinations; this latter a very human-centric metric related to human’s high value of historical truth.
@@
In Alan Turing’s famous test [7], it was suggested that intelligence is best judged by another intelligent system. This behaviourist approach defines intelligence not by its internal workings but by its external performance in tasks that require reasoning, problem solving, and communication. However, this perspective has been challenged by those who argue that true intelligence requires intrinsic understanding rather than mere imitation of intelligent behaviour. Douglas Hofstadter [8], for example, argued that defining intelligence in some other way than “that which gets the same meaning out of a sequence of symbols as we do” would support the idea of meaning being an inherent property.
@@
Some scholars argue that intelligence can be objectively defined through tests that evaluate specific computational abilities essential to demonstrate intelligent behaviour, rather than trying to define intelligence itself in absolute terms [2, 6, 3, 4]. This perspective shifts the focus from an abstract or philosophical definition to a practical, measurable framework assessing an entity’s capacity for problem-solving, pattern recognition, and adaptive learning within a structured system. In this regard, Gregory Chaitin [9] proposed that formal definitions of intelligence and measures of its various components, should be developed in the context of algorithmic complexity in application to AI. In the same vein, Solomonoff also explicitly proposed assessing AI capabilities using algorithmic probability [10] and is credited with having solved AI with its optimal inference theory [11].
@@
Based on these ideas, some tests for machine, human, and non-human entities have been proposed [12, 13, 4]. A generally accepted approach is that intelligence may be fundamentally linked to compression [5]–the ability to represent complex data in a simpler form while retaining meaning. This suggests that intelligence involves identifying patterns, making predictions, and generating concise explanations for observed phenomena. Such an approach provides a unified framework for understanding both human and artificial intelligence, moving beyond traditional tests and philosophical debates to a measurable and practical foundation.
@@
Similarly to a test proposed in [14], a benchmark designed to evaluate conceptual understanding in machine learning models was proposed [15] consisting of a diverse set of tasks that indirectly assess a model’s capacity for abstraction, requiring it to generalise beyond memorisation. These tasks challenge models to reason both interpolatively–by making sense of patterns within observed data–and extrapolatively–by extending learned principles to novel scenarios. While interesting and a first approach, the test lacked robust foundations of algorithmic complexity nor they were applied to frontier models.
@@
At recent public events, speaking about the foundations of AI and AGI, some AI-industry leaders have drawn strong parallels between algorithmic complexity, data compression, and AI [16, 17] making the connection between LLMs, algorithmic complexity and data compression more explicit, even calling it fundamental for general and super intelligence, artificial or natural. One idea expressed by Sutskever [16], is that Stochastic Gradient Descent (SGD), a main iterative optimisation algorithm for optimising an objective function used to train models in machine learning (ML) and artificial intelligence (AI), is a practical approximation to finding a computer program that compresses the encoding data in the search space and performs a type of ‘Kolmogorov search’ to find an implicit small computer program embedded in the weights of a ‘soft computer’ or a neural network such as a large Transformer. In a previous work, we successfully explored some of these ideas, proving that we can perform this search on non-differentiable spaces using metrics purely based on algorithmic complexity to search for those programs in model space, making the previously considered fundamental requirement of differentiality redundant [18]. Encoders are effectively lossy compression heuristics and therefore deeply connected to algorithmic complexity via compression.
@@
Building on our previous work reporting applications to various fields ranging from cell and molecular biology to genetics [19, 20] to biosignatures to animal and human behaviour [2, 3, 4], here we introduce a quantitative test for AGI and ASI with an application to LLMs fully framed in terms of the principles and foundations of Algorithmic Information Theory (AIT) [21, 22, 23, 24, 25, 26]. It is related to tests such as the ARC challenge [27], but is systematic, potentially more objective (since it does not pick specific test cases) and agnostic. We will illustrate the test in application to binary and integer sequences, but it is in no way limited to binary, integer, or even sequences for that matter, so as to avoid a metric that may become the target and cease to be useful. The new test is independent of, though connected to, the theory of mind and human intelligence, as demonstrated in the randomness perception and generation tests [2]. We will argue that an intelligent agent’s ability to find patterns (compression) is directly related to its ability to anticipate future events (planning and prediction), qualities that have recently been strongly associated with AI, AGI and ASI [28, 29].
@@ 2 Intelligence and Compression
This effective recursive decompression process not only reconstructs or reassembles the original explanandum but it can produce a continuation of it based on the continuation of the optimal recursive compressed features in reverse, producing a simulation that acts as a prediction on which a future action can be modelled. This amounts to the process of planning, as the outcome can be compared and adjusted by iterating over the recursive process, comparing the output against any evolving ground truth in a continuous learning process. This iterative updating process is the most optimal in the Bayesian sense [31, 32].
By proposing a formal and more objective definition of intelligence and based on our previous work on computational irreducibility and unpredictability [33], we propose a test for (Super)intelligence based on Algorithmic Information Theory (AIT) [34] specifically testing on features lately strongly associated with intelligence in the context of discussions of Artificial General Intelligence (AGI)  [28, 29, 35, 36], such as model abstraction, generalisation, and planning. Here we will argue that all or most of these features are related to just three, therefore one feature measured by three methods:
  • (Recursive) Compression and (recursive) decompression: seen as the abstraction of main features (or feature selection) that can be simulated in reverse (decompression);
  • Prediction: formally established by AIT as equivalent to compression by way of simulation [37, 38, 39] through the concept of algorithmic randomness and martingales (betting strategies) [40, 41, 42] (see Section Equivalence between compression and prediction via Martingales);
  • Regression: also part of AIT in the form of universal (Solomonoff) induction [10, 11, 23] (see also Section 1).
@@
3 Assessing the capabilities of frontier models and Large Language Models
@@
Other researchers, following a more abstract and formal approach, incline to the view that a test of intelligence in LLMs, which could imply comprehension, understanding, and prediction, might rely on exposing and training LLMs on complexity and not merely on intelligent data sets, and testing how well the LLMs could apply learned knowledge to unrelated but complex tasks (like predicting the next chess move) and reasoning tasks. They claim that information at the ‘edge of chaos’, a state between order and randomness, is more likely to help LLMs manifest intelligence  [55] as an emergent property. Suspicions that current AI is mimicking intelligence rather than displaying it have been reported and substantiated before [56, 54, 57]; therefore, proposing a test that can adequately address this issue is very relevant.
@@ 4 The SuperARC testing framework
This section introduces the key concepts required to define intelligence as considered in this work. Subsequently, we propose a general testing framework, referred to as SuperARC.
@@ 4.1 Foundations and Principles of Complexity Related to Intelligence
Intelligence is the ability to create a computable model that effectively (as losslessly as possible) explains any given data, where greater intelligence corresponds to more compact models.
In order to further develop the definition above, we use AIT to objectively describe what a compact model is, which is presented in the next sections.
4.1.1 Algorithmic Complexity
Algorithmic complexity, also referred to as Kolmogorov or Solomonoff-Kolmogorov-Chaitin complexity, is a measure of the complexity of a string of data or an object. The algorithmic complexity $K(\sigma)$ of a finite string $\sigma$ is the length of the shortest binary program (on a fixed universal Turing machine) that outputs $\sigma$. A string $\sigma$ is compressible if $K(\sigma)<|\sigma|$, where $|\sigma|$ is the length of $\sigma$. More complex objects require longer descriptions, while simpler, more regular objects can be described by shorter programs [21, 22, 58, 59].
Algorithmic complexity plays a crucial role in data compression, but goes well beyond compression. Consider, for example, a sequence of integers. The ability to compress such a sequence effectively is often taken as an indicator of understanding a model that is capable of generating the sequence, and one does not need to take the minimum requirement to the limit to find short plausible explanations. These explanations are mechanistic in nature as they can be built step-by-step by the universal constructor. The universal constructor is simply another computer program equivalent to a Turing machine (though not necessarily exactly a Turing machine). Solomonoff’s Theory of Inductive Inference proves that prediction and compression are tightly linked via universal induction. Solomonoff [23] also laid the foundation for Algorithmic probability, which is a universally optimal probability measure that a string is generated by a random program fed into a universal constructor or computer program (see Sup. Inf.).
@@ 4.1.2 Algorithmic Randomness and Intelligence
If a sequence $x$ can be represented by a shorter program $p$, the shorter program captures the regularities in $x$. In this sense, the program can be used to generate or predict future segments of the sequence, based on the learned regularities. Thus, the ability to compress is directly tied to the ability to predict future patterns.
In practical terms, compression algorithms like ZIP or LZW attempt to reduce the size of data by identifying recurring statistical patterns. If an AI system like ChatGPT can generate a concise and generalisable program to reproduce a sequence, it shows that the model has ‘compressed’ the information by finding underlying symbolic patterns. The latter is more powerful because it can continue generating data while statistical pattern matching does not. Pattern matching can only be descriptive, but symbolic regression and program synthesis can be prescriptive.
A key aspect of algorithmic complexity is this deeper relationship with randomness, in comparison to statistical randomness, defined as a lack of statistical patterns. A sequence is considered algorithmically random if its shortest description is essentially the sequence itself, i.e., no shorter program exists to generate it (i.e., it can at best be described as a program of the type ‘print($x$)’). Mathematically, a string $x$ is random if $K(x)\approx|x|$, where $|x|$ is the length of the string in bits. In this case, $x$ is incompressible because no smaller program can produce it, which contrasts with highly structured or predictable data, where $K(x)\ll|x|$. When a statistical compression algorithm such as ZIP or LZW compresses $x$, it is a sufficient proof of non-randomness. However, if it does not compress $x$, it will keep it about the same size and will not be a proof of non-randomness because there may be a program that statistical compression is unable to produce.
In the theory of algorithmic randomness, Schnorr and Levin independently established a profound connection between prediction and compression [41, 42, 38, 39]. They proved that a sequence is algorithmically random if and only if no computable betting strategy (martingale) can succeed on it. This result demonstrated that the ability to compress a sequence is equivalent to the inability to predict its future bits using any effective method. Proof of this equivalence using martingales is provided in the Supplementary Material.
A sequence is algorithmically random (incompressible) if and only if no computable martingale succeeds on it. This establishes the equivalence between the inability to compress a sequence and the impossibility of predicting its future bits using any computable betting strategy.
This equivalence highlights the deep interplay between randomness, prediction, and compression in the context of algorithmic information theory, as established by Schnorr and Levin using martingales [60, 41, 38, 42].
@@
Chaitin showed that a random string cannot be significantly compressed [22], implying that intelligence (as seen in systems that can compress data) involves recognising non-random patterns in data. Universal Predictors (like those based on Levin’s universal search [31] (Sup Inf.) or universal induction [23]) use algorithmic complexity [21] to model the most likely future based on past data, effectively capturing the link between compression and prediction.
Large Language Models (LLMs) can be thought of as word time series predictors based on short- and long-range correlations that compress data from their very large training sets based on text repositories mostly available online, and captured in a much smaller object such as a giant matrix, whose numerical entries can partially and lossily reconstruct the training dataset. Whether they build a compressed version that can amount to a level of understanding or comprehension is what this work and test sets out to help assess and determine, based on the correct algorithmic framework.

### v2

# SuperARC: An Agnostic Test for Narrow, General, and Super Intelligence Based On the Principles of Recursive Compression and Algorithmic Probability

@@ Abstract
We introduce an open-ended test grounded in algorithmic probability that can avoid benchmark contamination in the quantitative evaluation of frontier models in the context of their Artificial General Intelligence (AGI) and Superintelligence (ASI) claims. Unlike other tests, this test does not rely on statistical compression methods (such as GZIP or LZW), which are more closely related to Shannon entropy than to Kolmogorov complexity and are not able to test beyond simple pattern matching. The test challenges aspects of AI, in particular LLMs, related to features of intelligence of fundamental nature such as synthesis and model creation in the context of inverse problems (generating new knowledge from observation). We argue that metrics based on model abstraction and abduction (optimal Bayesian ‘inference’) for predictive ‘planning’ can provide a robust framework for testing intelligence, including natural intelligence (human and animal), narrow AI, AGI, and ASI. We found that LLM model versions tend to be fragile and incremental as a result of memorisation only with progress likely driven by the size of training data. The results were compared with a hybrid neurosymbolic approach that theoretically guarantees universal intelligence based on the principles of algorithmic probability and Kolmogorov complexity. The method outperforms LLMs in a proof-of-concept on short binary sequences. We prove that compression is equivalent and directly proportional to a system’s predictive power and vice versa. That is, if a system can better predict it can better compress, and if it can better compress, then it can better predict. Our findings strengthen the suspicion regarding the fundamental limitations of LLMs, exposing them as systems optimised for the perception of mastery over human language.
Keywords: ARC tests, prediction, compression, program synthesis, inverse problems, symbolic regression, comprehension, Superintelligence, Generative AI, symbolic computation, hybrid computation, Neurosymbolic computation.
@@ 1 Introduction
We are heavily biased to believe that the way humans think and act represents the acme of intelligence, even in instances where we may be limited, or flawed or irrational, or engaged in narrowly specific human (and often mundane) activities like chatting or washing dishes.
@@
One of the greatest realisations from the impressive apparent performance of Large Language Models (LLMs) is that language and other areas of human intellect may be overrated and are more dependent than we thought on memorisation and statistical pattern matching than critical thinking or other features of general intelligence.
One of the first metrics for intelligence was introduced by Charles Spearman in 1904 [1]. He proposed specific tests called ‘s’ that would each contribute to a general intelligence test under the name ‘g’, representing the common cognitive ability underlying performance in various mental tasks. Specific intelligences that contribute to the estimation of the g factor are verbal comprehension, perceptual reasoning, working memory, processing speed, quantitative reasoning, abstract reasoning, spatial ability, memory retrieval, auditory processing, and fluid reasoning. Some LLM benchmarks test for different factors, with several benchmarks based on correct answers versus hallucinations; some of which are very human-centric metrics related to human’s biological peculiarities and shared history.
@@
In Alan Turing’s famous test [7], it was described how an intelligent system could judge the intelligence of another system. This behaviourist approach defines intelligence not by its internal workings but by its external performance in tasks that require reasoning, problem solving, and communication. However, this perspective has been challenged by those who argue that true intelligence requires intrinsic understanding rather than mere imitation of intelligent behaviour. Douglas Hofstadter [8], for example, argued that defining intelligence in some other way than “that which gets the same meaning out of a sequence of symbols as we do” would support the idea of meaning being an inherent property.
@@
Some scholars argue that intelligence can be objectively defined through tests that evaluate specific computational abilities essential to demonstrate intelligent behaviour, rather than trying to define intelligence itself in absolute terms [2, 6, 3, 4]. This perspective shifts the focus from an abstract or philosophical definition to a practical, measurable framework assessing an entity’s capacity for problem-solving, pattern recognition, and adaptive learning within a structured system. In this regard, Gregory Chaitin [9] proposed that formal definitions of intelligence and measures of its various components should be developed in the context of algorithmic complexity in its application to AI. In the same vein, Solomonoff also explicitly proposed assessing AI capabilities using algorithmic probability [10] and is credited with having solved AI with its optimal inference theory [11].
@@
Based on these ideas, some tests for machine, human, and non-human entities have been proposed [12, 13, 4]. A generally accepted approach is that intelligence may be fundamentally linked to compression [5]—i.e., the ability to represent complex data in a simpler form while trying to lose the least information as possible. This suggests that intelligence involves identifying patterns, making predictions, and generating concise explanations for observed phenomena. Such an approach provides a unified framework for understanding both human and artificial intelligence, moving beyond traditional tests and philosophical debates to a measurable and practical foundation.
@@
Similarly to a test proposed in [14], a benchmark designed to evaluate conceptual understanding in machine learning models was proposed [15] consisting of a diverse set of tasks that indirectly assess a model’s capacity for abstraction, requiring it to generalise beyond memorisation. These tasks challenge models to reason both interpolatively (by making sense of patterns within observed data) and extrapolatively (by extending learnt principles to novel scenarios). Although interesting and a first approach, the test lacked robust foundations of algorithmic complexity, nor were they applied to frontier models.
@@
At recent public events, speaking about the foundations of AI and AGI, some leaders in the AI industry have drawn strong parallels between algorithmic complexity, data compression, and AI [16, 17]. Although these terminologies, such as AGI and ASI, are currently loosely defined in the scientific literature, these claims and the current understanding make the connection between LLMs (or any other generative AI), algorithmic complexity, and data compression clearer and more explicit, even calling it fundamental for general and super intelligence, artificial or natural. One idea expressed by Sutskever [16], is that Stochastic Gradient Descent (SGD), a main iterative optimisation algorithm for optimising an objective function used to train models in machine learning (ML) and artificial intelligence (AI), is a practical approximation to finding a computer program that compresses the encoding data in the search space and performs a type of ‘Kolmogorov search’ to find an implicit small computer program embedded in the weights of a ‘soft computer’ or a neural network such as a large Transformer. In a previous work, we successfully explored some of these ideas, proving that we can perform this search on non-differentiable spaces using metrics purely based on algorithmic complexity to search for those programs in model space, making the previously considered fundamental requirement of differentiality redundant [18]. Encoders are effectively (lossy or lossless) compression heuristics and, therefore, deeply connected to algorithmic complexity via compression. Similar ideas are also in evidence in Schmidhuber’s Gödel machines [19] work and Hutter’s AIXI [11] based on Levin’s search [20] and the principles of Algorithmic Probability [21].
@@
Building on our previous work reporting applications to various fields ranging from cell and molecular biology to genetics [22, 23] to biosignatures to animal and human behaviour [2, 3, 4], here we introduce a quantitative test for any AI model that aims at universal optimization and problem-agnostic capabilities (therefore, a test for what can be understood as AGI and ASI) with an application to LLMs fully framed in terms of the principles and foundations of Algorithmic Information Theory (AIT) [24, 25, 21, 26, 27, 28]. It is related to tests such as the ARC challenge [29], but is systematic, potentially more objective (since it does not pick specific test cases) and agnostic. We will illustrate the test in application to binary and integer sequences, but it is in no way limited to binary, integer, or even sequences for that matter, so as to avoid a metric that may become the target and cease to be useful. The new test is independent of, though connected to, the theory of mind and human intelligence, as demonstrated in the randomness perception and generation tests [2]. We will argue that an intelligent agent’s ability to find patterns (compression) is directly related to its ability to anticipate future events (planning and prediction), qualities that have recently been strongly associated with AI, AGI and ASI [30, 31].
@@ 2 Intelligence and Compression
This effective recursive decompression process not only reconstructs or reassembles the original explanandum but it can produce a continuation of it based on the continuation of the optimal recursive compressed features in reverse, producing a simulation that acts as a prediction on which a future action can be modelled. This amounts to the process of planning, as the outcome can be compared and adjusted by iterating over the recursive process, comparing the output against any evolving ground truth in a continuous learning process. This iterative update process is the most optimal in the Bayesian sense [20, 33].
By proposing a formal and more objective definition of intelligence and based on our previous work on computational irreducibility and unpredictability [34], we propose a test for (Super)intelligence based on Algorithmic Information Theory (AIT) [35] specifically testing recently strongly associated features with intelligence in the context of discussions of Artificial General Intelligence (AGI)  [30, 31, 36, 37], such as model abstraction, generalisation, and planning. Here we will argue that all or most of these features are related to just three, therefore, one feature measured by three methods:
  • Recursive Compression and recursive decompression: seen as the abstraction of main features (or feature selection) that can be simulated in reverse (decompression) and in contrast to simple statistical pattern-matching or statistical compression;
  • Symbolic Regression and Prediction: formally established by AIT as equivalent to compression by way of optimal simulation [38, 39, 40] through the concept of algorithmic randomness and martingales (betting strategies) [41, 42, 43] (see Section 10.3); or universal (Solomonoff) induction [10, 11, 21] (see also pseudocode 1).
@@
3 Assessing the capabilities of frontier models and LLMs
@@
Other researchers, following a more abstract and formal approach, incline to the view that a test of intelligence in LLMs, which could imply comprehension, understanding, and prediction, might rely on exposing and training LLMs on complexity and not merely on intelligent data sets, and testing how well the LLMs could apply learnt knowledge to unrelated but complex tasks (like predicting the next chess move) and reasoning tasks. They claim that information at the ‘edge of chaos’, a state between order and randomness, is more likely to help LLMs manifest intelligence  [55] as an emergent property. Suspicions that current AI is mimicking intelligence rather than displaying it have been reported and substantiated before [56, 54, 57]; therefore, proposing a test that can adequately address this issue is very relevant.
@@ 4 The SuperARC testing framework
We propose a general testing framework, referred to as SuperARC.
@@ 4.1 Foundations and Principles of Complexity Related to Intelligence
Intelligence is the ability to create a computable model that effectively (as losslessly as possible) explains any given data, where greater intelligence corresponds to performing an optimal prediction (abduction) from compact model representations.
The technical framework of the definition above is the theory of algorithmic information which is a generalisation of classical Shannon information theory and the accepted mathematical definition that tells apart randomness from non-randomness (mechanical causality) able to objectively describe and quantify what a compact model is and what optimal prediction (induction.abduction) means.
4.1.1 Algorithmic Information Theory (AIT)
Algorithmic complexity, also referred to as Kolmogorov or Solomonoff-Kolmogorov-Chaitin complexity, is at the centre of AIT and is a measure of the complexity of a string of data or an object. The algorithmic complexity $K(\sigma)$ of a finite string $\sigma$ is the length of the shortest binary program (on a fixed universal Turing machine) that outputs $\sigma$. A string $\sigma$ is compressible if $K(\sigma)<|\sigma|$, where $|\sigma|$ is the length of $\sigma$. More complex objects require longer descriptions, while simpler, more regular objects can be described by shorter programs [24, 25, 58, 59].
Algorithmic complexity goes beyond strings, beyond binary and beyond computer programs. It only uses this language as a technicality given the fundamental nature of strings, binary language, and computer programs. For example, as proven by Shannon any discrete data can be transformed to binary without loss of information, any computable description and rule can be described as a computer program under the Church-Turing thesis, which underlies all science as it presumes and assumes that world phenomena can objectively be described in a form in which science can process or deal this process and data with (e.g. equations, computer simulations). These computer programs are also not restricted to deal with strings only, just as computers deal with images, vectors, tensors, sounds, video or anything else.
Algorithmic complexity is therefore a concept of fundamental nature in science and even if it also plays a crucial role in data compression, but goes well beyond compression [5]. Science itself can fundamentally be seen as compressing, as the process of producing ever more compact representations of the physical world into rules, equations, and scientific models. that provide ever greater predicting power.
For illustration purposes and without loss of generality, let us consider a sequence of integers. The ability to compress such a sequence effectively is often taken as an indicator of understanding a model that is capable of generating the sequence, and one does not need to take the minimum requirement to the limit to find short plausible explanations. These explanations are mechanistic in nature as they can be built step-by-step by the universal constructor. The universal constructor is simply another computer program equivalent to a Turing machine (though not necessarily exactly a Turing machine). Solomonoff’s Theory of Inductive Inference proves that prediction and compression are tightly linked via universal induction (or abduction). Solomonoff [21] also laid the foundation for Algorithmic probability, which is a universally optimal probability measure in which a string is generated by a random program fed into a universal constructor or computer program (see Sup. Inf.).
@@ 4.1.2 Algorithmic Randomness and Intelligence
If a sequence $x$ can be represented by a shorter program $p$, the shorter program captures the regularities in $x$. In this sense, the program can be used to generate or predict future segments of the sequence, based on the learnt regularities. Thus, the ability to compress is directly tied to the ability to predict future patterns.
In practical terms, compression algorithms like ZIP or LZW attempt to reduce the size of the data by identifying recurring statistical patterns. If an AI system like ChatGPT can generate a concise and generalisable program to reproduce a sequence, it shows that the model has ‘compressed’ the information by finding underlying symbolic patterns. The latter is more powerful because it can continue generating data while statistical pattern matching does not. Pattern matching can only be descriptive, but symbolic regression and program synthesis can be prescriptive.
A key aspect of algorithmic complexity is this deeper relationship with randomness, in comparison to statistical randomness, defined as a lack of statistical patterns. A sequence is considered algorithmically random if its shortest description is essentially the sequence itself, i.e., no shorter program exists to generate it (i.e., it can at best be described as a program of the type ‘print($x$)’). Mathematically, a string $x$ is random if $K(x)\geq|x|-\mathbf{O}(1)$, where $|x|$ is the length of the string in bits. In this case, $x$ is incompressible because no smaller program can produce it, which contrasts with highly structured or predictable data, where $K(x)\ll|x|$. When a statistical compression algorithm such as ZIP or LZW compresses $x$, it is a sufficient proof of non-randomness. However, if it does not compress $x$, it will keep it about the same size and will not be a proof of non-randomness because there may be a program that statistical compression is unable to produce.
The theory of algorithmic randomness, established a profound connection between prediction and compression [42, 43, 39, 40]. They proved that a sequence is algorithmically random if and only if no computable betting strategy (martingale) can succeed on it. This result demonstrated that the ability to compress a sequence is equivalent to the inability to predict its future bits using any effective method. Proof of this equivalence using martingales is provided in the Sup. Inf.. A random string cannot be significantly compressed [25], implying that intelligence (as seen in systems that can compress data) involves recognising non-random patterns in data. Therefore, it is equivalent to say that a sequence is algorithmically random (incompressible) iff no computable martingale succeeds on it, establishing the equivalence between the inability to compress a sequence and the impossibility of predicting its future bits using any computable betting strategy. This also highlights the deep interplay between randomness, prediction, and compression in the context of algorithmic information theory.
@@
Universal Predictors (like those based on Levin’s universal search [20] (Sup. Inf.) or universal induction [21]) use algorithmic complexity [24] to model the most likely future based on past data, effectively capturing the link between compression and prediction.
Large Language Models (LLMs) can be thought of as word (token) time series predictors based on short- and long-range correlations that compress data from their very large training sets based on text repositories mostly available online, and captured in a much smaller object such as a giant matrix, whose numerical entries can partially and lossy reconstruct the training dataset. Whether they build a compressed version that can amount to a level of understanding or comprehension is what this work (and test) sets out to help assess and determine, based on the correct algorithmic framework.